The road at a circular turn of radius 10 m is bankred by an angle of 10. With what speed should a vehicle move on the turn so that the normal contact force is able to provide the necessary centripetal force ?

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
If \nuis the correct speed,

\tan \theta = \frac{v^{2}}{rg}

or,

\nu = \sqrt{rg\tan \theta } = \sqrt{(10m)(9.8 m/s^{2})(tan 10^{\circ})} = 4.2 m/s.